### v1

@@ 6.1 Performance analysis on the IRP
Classical benchmark instances for the IRP do not permit stock-out. To accommodate this setting, we omit negative $\hat{I}_{i}^{t}$ in the DP transition processes described in Section 4 while retaining the rest of the algorithm unchanged. We compare the performance of our algorithm, HGS-IRP, against the most recent heuristics in the literature as well as the best-performing method from the DIMACS 2022 Implementation Challenge on Vehicle Routing, which included a dedicated track on the IRP. All benchmark instances and reference solutions used in our experiments are sourced from https://or-brescia.unibs.it/instances and http://dimacs.rutgers.edu/programs/challenge/vrp/irp/, and are also provided in the Git repository associated with this project. HGS-IRP is evaluated against three leading methods: the unified decomposition matheuristic proposed by Chitsaz et al. (2019) (denoted CCJ), the kernel search heuristic of Archetti et al. (2021) (KS), and the Branch-and-Heuristic approach developed by Skålnes et al. (2023) (BCHeu), which achieved first place in the DIMACS competition. This selection provides a comprehensive basis for performance comparison.
Table 1: Average results on benchmark IRP datasets.
@@
$H$ $C$ $n$ Sol Gap(%) T(s) Sol Gap(%) T(s) Sol Gap(%) T(s) Avg Gap(%) T(s) Best BKS
3 HC 5 2855.05 3.39 1.7 2761.47 0.00 5.9 2761.47 0.00 7200.0 2761.47 0.00 60.3 2761.47 2761.47
10 5013.81 3.24 4.3 4856.89 0.01 91.3 4856.50 0.00 7200.0 4856.50 0.00 104.3 4856.50 4856.50
15 5620.86 1.63 11.3 5534.01 0.06 114.9 5530.87 0.00 7200.0 5530.88 0.00 119.6 5530.87 5530.87
20 7191.44 1.39 25.3 7105.34 0.18 202.6 7093.86 0.02 7200.0 7092.57 0.00 127.0 7092.57 7092.57
25 8501.63 0.86 37.6 8438.19 0.10 324.3 8429.51 0.00 7200.0 8430.00 0.01 153.9 8429.50 8429.42
30 9699.83 0.40 64.1 9721.92 0.63 504.8 9665.34 0.04 7200.0 9660.65 0.00 207.5 9660.57 9661.11
35 10210.21 0.39 85.2 10205.92 0.35 544.2 10172.89 0.02 7200.0 10171.15 0.01 259.5 10170.59 10170.59
40 10986.12 0.46 141.9 11012.14 0.70 614.7 10945.29 0.09 7200.0 10950.63 0.13 318.9 10935.93 10935.94
45 11998.55 0.27 179.9 12077.35 0.93 663.1 11999.87 0.28 7200.0 11976.60 0.09 313.1 11966.17 11966.14
50 13571.12 0.33 232.1 13692.82 1.23 1173.8 13546.34 0.15 7200.0 13528.67 0.02 453.9 13523.46 13526.27
Avg 8564.86 0.85 78.3 8540.60 0.56 424.0 8500.19 0.08 7200.0 8495.91 0.03 211.8 8492.76 8493.09
LC 5 2180.27 4.81 1.7 2080.31 0.00 6.2 2080.31 0.00 7200.0 2080.32 0.00 60.4 2080.31 2080.31
10 3221.39 5.51 4.2 3053.18 0.00 89.3 3053.14 0.00 7200.0 3053.13 0.00 99.4 3053.13 3053.13
15 3200.77 2.66 10.3 3121.36 0.11 118.5 3118.09 0.01 7200.0 3117.85 0.00 107.6 3117.85 3117.85
20 3771.04 2.27 20.5 3702.66 0.41 225.8 3687.81 0.01 7200.0 3687.49 0.00 128.6 3687.48 3687.48
25 3994.91 1.49 60.6 3951.99 0.40 337.7 3937.48 0.03 7200.0 3938.14 0.04 149.6 3936.56 3936.37
30 3989.68 1.07 44.1 3978.59 0.78 482.1 3953.57 0.15 7200.0 3947.76 0.00 192.3 3947.63 3947.63
35 4237.04 0.94 58.1 4224.87 0.65 556.9 4198.92 0.03 7200.0 4197.83 0.01 207.2 4197.62 4197.58
40 4397.24 1.89 94.4 4377.20 1.43 619.9 4328.67 0.30 7200.0 4322.94 0.17 264.5 4315.28 4315.63
45 4517.73 1.03 107.4 4576.23 2.33 698.3 4491.29 0.44 7200.0 4473.40 0.04 274.6 4471.83 4471.83
50 5296.46 1.16 152.7 5445.02 4.00 1098.2 5253.17 0.34 7200.0 5239.09 0.07 365.5 5232.78 5235.53
Avg 3880.65 2.01 55.4 3851.14 1.23 423.3 3810.24 0.16 7200.0 3805.79 0.04 185.0 3798.08 3804.33
6 HC 5 7371.68 0.00 95.6 7371.44 0.00 7200.0 7090.34 -3.81 394.3 7090.08 7371.44
10 11760.57 7.54 13.1 10954.19 0.17 134.1 10936.89 0.01 7200.0 10940.78 0.05 828.9 10935.72 10935.67
15 13797.25 5.31 77.7 13156.04 0.42 198.5 13110.17 0.07 7200.0 13117.52 0.12 1042.8 13105.75 13101.33
20 16979.09 4.03 71.5 16420.79 0.61 503.7 16334.12 0.08 7200.0 16335.86 0.09 1358.5 16320.20 16320.60
25 19081.64 3.97 135.5 18498.31 0.79 694.3 18395.07 0.23 7200.0 18376.31 0.13 1590.9 18356.10 18353.33
30 21314.31 2.41 221.3 21034.43 1.06 687.2 20866.21 0.25 7200.0 20836.78 0.11 1766.2 20812.69 20813.61
35 22671.70 0.29 7200.0 22640.50 0.15 1888.2 22609.81 22606.89
40 24292.25 0.63 7200.0 23451.28 -2.86 1923.8 23400.59 24141.23
45 26722.37 0.21 7200.0 25920.28 -2.80 2007.8 25880.87 26667.58
50 30467.00 0.15 7200.0 29567.18 -2.81 2201.4 29515.75 30420.51
Avg 19116.72 0.23 7200.0 18827.68 -1.16 1500.27 18802.75 19073.22
LC 5 5794.88 5.78 5.8 5478.02 0.00 88.7 5477.99 0.00 7200.0 5478.03 0.00 428.7 5477.99 5477.99
10 8213.51 9.15 11.0 7548.02 0.31 130.0 7525.02 0.00 7200.0 7526.81 0.03 725.4 7525.61 7524.83
15 8469.11 7.79 27.2 7891.16 0.44 220.5 7863.14 0.08 7200.0 7869.35 0.16 897.9 7858.85 7856.69
20 10123.49 6.94 59.1 9558.74 0.97 641.7 9506.25 0.42 7200.0 9488.26 0.23 1260.5 9476.98 9466.51
25 10648.25 5.86 108.4 10227.55 1.68 723.7 10104.99 0.46 7200.0 10088.21 0.29 1525.1 10064.89 10058.85
30 10394.63 4.84 153.9 10222.89 3.10 619.1 9971.14 0.57 7200.0 9933.83 0.19 1735.6 9901.21 9915.06
35 10915.93 0.79 7200.0 10843.66 0.12 1925.4 10817.43 10830.43
40 11002.22 0.27 7200.0 10991.20 0.17 1985.1 10951.25 10972.65
45 11579.88 0.46 7200.0 11544.00 0.15 2025.8 11510.08 11527.10
50 13511.62 0.26 7200.0 13505.18 0.21 2262.0 13454.90 13476.69
Avg 9745.82 0.36 7200.0 9726.85 0.17 1477.2 9703.92 9710.68
6 LC 50 14336.56 4.96 394.5 14455.92 5.84 2589.5 13694.50 0.26 7200.0 13670.93 0.09 4512.1 13619.39 13658.58
100 17630.85 4.50 1823.2 18421.73 9.18 2112.0 16890.54 0.11 7200.0 16805.40 -0.40 6391.5 16736.78 16872.38
200 25142.35 4.12 8310.1 27786.21 15.07 828.3 24146.93 0.00 7200.0 24216.63 0.29 7140.9 24063.84 24146.93
Avg 19036.59 4.45 3509.3 20221.29 10.95 1843.3 18243.99 0.10 7200.0 18230.99 0.03 6014.8 18140.00 18225.96
HC 50 31283.24 2.15 642.5 31412.44 2.57 3081.1 30663.33 0.13 7200.0 30649.56 0.08 4323.8 30581.14 30624.92
100 52530.23 1.30 2968.6 53591.81 3.35 2360.4 51865.07 0.02 7200.0 51888.67 0.06 5531.8 51740.12 51856.76
200 96076.10 0.82 16435.8 99299.77 4.20 1507.6 95298.01 0.00 7200.0 95576.34 0.29 6927.1 95358.71 95298.01
Avg 59963.19 1.19 6682.3 61434.67 3.67 2316.4 59275.47 0.03 7200.0 59371.52 0.19 5594.2 59226.66 59259.90
Avg Gap 2.84 2.66 0.17 -0.15
@@
Results are summarized in Table 1, which reports average values for instances sharing the same values of $n$, $H$, and $C$, on the benchmark instances, small and large (the latter corresponding to the last two blocks). For each method, we report the solution value (Sol), the percentage gap (Gap) relative to the best known solution (BKS, shown in the last column), and the computational time in seconds (T(s)). The gap is computed as $\left(\frac{\mathrm{Sol}}{\mathrm{BKS}}-1\right)\times 100\%$. To the best of our knowledge, the BKS values are compiled from all existing methods for the IRP under an ML-based inventory policy, including the best-performing solutions reported in the DIMACS 2022 Vehicle Routing Implementation Challenge. In contrast to these single-run results, we report the average results of our method over multiple runs, as this provides a more reliable performance estimator given the random nature of the algorithm. Specifically, we present the average solution value (Avg), average optimality gap (Gap), average CPU time (T(s)), and the best solution among the 10 runs (Best).
The proposed HGS-IRP algorithm exhibits strong performance in both solution quality and computational efficiency. It consistently produces solutions that are comparable to, or better than, those obtained by existing methods, all within reasonable computational times. Across all 1,040 tested instances, the maximum optimality gap relative to the BKS is only 0.29%, while the average gap is $-0.15\%$, both of which outperform all benchmark algorithms. In terms of computational time, HGS-IRP is particularly efficient on large-scale instances, and it also maintains strong solution quality on smaller instances. As the problem size increases, both in number of periods ($H$) and customers ($n$), the advantages of HGS-IRP become more pronounced. On the small 3-day dataset (400 instances), HGS-IRP identifies 37 new best-known solutions, matches the existing BKS in 320 instances, and performs slightly worse in only 43 instances. For the small 6-day dataset, it finds 150 new best solutions and matches the BKS in 83 instances. Most notably, on the large 6-day dataset, HGS-IRP finds new best solutions in 178 out of 240 instances, representing a substantial improvement over the current state-of-the-art in solving large-scale IRP problems.
@@
Experiments on SAHS instances. We further evaluate HGS-IRP on the challenging IRP instances recently introduced by Skålnes et al. (2024). Table 2 presents the average results across instances of varying sizes. Skålnes et al. (2024) propose three approaches, and we report the corresponding results: (1) Construction heuristic, (2) Matheuristic, (3) Branch-and-Cut. Only the Branch-and-Cut method provides lower bounds (LB), in contrast to the constructive and matheuristic methods, which are primarily designed to generate feasible solutions and, therefore, only provide upper bounds (UB). Finally, BKS reports the best of the feasible solutions found by the three approaches (the minimum value in UB).
@@
The computational results demonstrate that, out of a total of 270 benchmark instances, the best-known solution is improved in 191 instances and matched in an additional 25 instances. Notably, within the same 2-hour time limit, substantially better solutions are obtained—particularly on large-scale instances—yielding an average improvement of 24.01% over the best solutions previously reported by Skålnes et al. (2024). This average is computed across these 216 instances (i.e., 191 improved and 25 matched), where the improvement for each instance is defined as the relative gap reduction.
Table 2: Comparison with best-known UB on new challenging IRP dataset:
@@
$n$ UB Gap(%) T(s) UB Gap(%) T(s) LB UB Gap(%) T(s) Avg Gap(%) T(s) Best BKS
10 7502.74 40.08 0.0 5385.52 0.55 1748.8 5331.14 5356.03 0.00 5422.3 5369.76 0.26 4378.4 5356.81 5356.03
25 19431.74 40.37 0.0 15072.20 8.88 1800.0 13626.19 13842.88 0.00 7199.1 13886.24 0.31 7078.1 13819.07 13842.88
50 42969.72 6.84 0.1 - - 2400.0 28862.31 51389.61 27.78 7193.2 29945.03 -25.54 7390.6 29753.66 40218.66
100 86098.23 0.38 0.7 - - 2400.0 52307.74 172930.13 101.61 7193.8 58382.45 -31.94 7215.3 58121.23 85775.08
200 154146.38 0.00 6.7 - - 2400.0 76747.91 - - 7217.9 106730.50 -30.76 11139.3 105188.88 154146.38

### v2

@@ 6.1 Performance analysis on the IRP
Classical benchmark instances for the IRP do not permit stock-out. To accommodate this setting, we omit negative $\hat{I}_{i}^{t}$ in the DP transition processes described in Section 4 while retaining the rest of the algorithm unchanged. We compare the performance of our algorithm, HGS-IRP, against the most recent heuristics in the literature as well as the best-performing method from the DIMACS 2022 Implementation Challenge on Vehicle Routing, which included a dedicated track on the IRP. All benchmark instances and reference solutions used in our experiments are sourced from https://or-brescia.unibs.it/instances and http://dimacs.rutgers.edu/programs/challenge/vrp/irp/, and are also provided in the Git repository associated with this project. HGS-IRP is evaluated against three leading methods: the unified decomposition matheuristic proposed by chitsaz2019unified (denoted CCJ), the kernel search heuristic of archetti2021kernel (KS), and the Branch-and-Heuristic approach developed by skaalnes2023branch (BCHeu), which achieved first place in the DIMACS competition. This selection provides a comprehensive basis for performance comparison.
Table 1: Comparative analysis on medium-scale IRP instances
@@
$H$ $C$ $n$ Sol Gap(%) T(s) Sol Gap(%) T(s) Sol Gap(%) T(s) Avg $\mathrm{Gap}_{\mathrm{A}}$(%) T(s) Best $\mathrm{Gap}_{\mathrm{B}}$(%) BKS
3 HC 5 2855.05 3.39 1.7 2761.47 0.00 5.9 2761.47 0.00 7200.0 2761.47 0.00 60.3 2761.47 0.00 2761.47
10 5013.81 3.24 4.3 4856.89 0.01 91.3 4856.50 0.00 7200.0 4856.50 0.00 104.3 4856.50 0.00 4856.50
15 5620.86 1.63 11.3 5534.01 0.06 114.9 5530.87 0.00 7200.0 5530.88 0.00 119.6 5530.87 0.00 5530.87
20 7191.44 1.39 25.3 7105.34 0.18 202.6 7093.86 0.02 7200.0 7092.57 0.00 127.0 7092.57 0.00 7092.57
25 8501.63 0.86 37.6 8438.19 0.10 324.3 8429.51 0.00 7200.0 8430.00 0.01 153.9 8429.50 0.00 8429.42
30 9699.83 0.40 64.1 9721.92 0.63 504.8 9665.34 0.04 7200.0 9660.65 0.00 207.5 9660.57 -0.01 9661.11
35 10210.21 0.39 85.2 10205.92 0.35 544.2 10172.89 0.02 7200.0 10171.15 0.01 259.5 10170.59 0.00 10170.59
40 10986.12 0.46 141.9 11012.14 0.70 614.7 10945.29 0.09 7200.0 10950.63 0.13 318.9 10935.93 0.00 10935.94
45 11998.55 0.27 179.9 12077.35 0.93 663.1 11999.87 0.28 7200.0 11976.60 0.09 313.1 11966.17 0.00 11966.14
50 13571.12 0.33 232.1 13692.82 1.23 1173.8 13546.34 0.15 7200.0 13528.67 0.02 453.9 13523.46 -0.02 13526.27
Avg 8564.86 0.85 78.3 8540.60 0.56 424.0 8500.19 0.08 7200.0 8495.91 0.03 211.8 8492.76 0.00 8493.09
LC 5 2180.27 4.81 1.7 2080.31 0.00 6.2 2080.31 0.00 7200.0 2080.32 0.00 60.4 2080.31 0.00 2080.31
10 3221.39 5.51 4.2 3053.18 0.00 89.3 3053.14 0.00 7200.0 3053.13 0.00 99.4 3053.13 0.00 3053.13
15 3200.77 2.66 10.3 3121.36 0.11 118.5 3118.09 0.01 7200.0 3117.85 0.00 107.6 3117.85 0.00 3117.85
20 3771.04 2.27 20.5 3702.66 0.41 225.8 3687.81 0.01 7200.0 3687.49 0.00 128.6 3687.48 0.00 3687.48
25 3994.91 1.49 60.6 3951.99 0.40 337.7 3937.48 0.03 7200.0 3938.14 0.04 149.6 3936.56 0.00 3936.37
30 3989.68 1.07 44.1 3978.59 0.78 482.1 3953.57 0.15 7200.0 3947.76 0.00 192.3 3947.63 0.00 3947.63
35 4237.04 0.94 58.1 4224.87 0.65 556.9 4198.92 0.03 7200.0 4197.83 0.01 207.2 4197.62 0.00 4197.58
40 4397.24 1.89 94.4 4377.20 1.43 619.9 4328.67 0.30 7200.0 4322.94 0.17 264.5 4315.28 -0.01 4315.63
45 4517.73 1.03 107.4 4576.23 2.33 698.3 4491.29 0.44 7200.0 4473.40 0.04 274.6 4471.83 0.00 4471.83
50 5296.46 1.16 152.7 5445.02 4.00 1098.2 5253.17 0.34 7200.0 5239.09 0.07 365.5 5232.78 -0.05 5235.53
Avg 3880.65 2.01 55.4 3851.14 1.23 423.3 3810.24 0.16 7200.0 3805.79 0.04 185.0 3798.08 -0.16 3804.33
6 HC 5 7371.68 0.00 95.6 7371.44 0.00 7200.0 7090.34 -3.81 394.3 7090.08 -3.82 7371.44
10 11760.57 7.54 13.1 10954.19 0.17 134.1 10936.89 0.01 7200.0 10940.78 0.05 828.9 10935.72 0.00 10935.67
15 13797.25 5.31 77.7 13156.04 0.42 198.5 13110.17 0.07 7200.0 13117.52 0.12 1042.8 13105.75 0.03 13101.33
20 16979.09 4.03 71.5 16420.79 0.61 503.7 16334.12 0.08 7200.0 16335.86 0.09 1358.5 16320.20 0.00 16320.60
25 19081.64 3.97 135.5 18498.31 0.79 694.3 18395.07 0.23 7200.0 18376.31 0.13 1590.9 18356.10 0.02 18353.33
30 21314.31 2.41 221.3 21034.43 1.06 687.2 20866.21 0.25 7200.0 20836.78 0.11 1766.2 20812.69 0.00 20813.61
35 22671.70 0.29 7200.0 22640.50 0.15 1888.2 22609.81 0.01 22606.89
40 24292.25 0.63 7200.0 23451.28 -2.86 1923.8 23400.59 -3.07 24141.23
45 26722.37 0.21 7200.0 25920.28 -2.80 2007.8 25880.87 -2.95 26667.58
50 30467.00 0.15 7200.0 29567.18 -2.81 2201.4 29515.75 -2.97 30420.51
Avg 19116.72 0.23 7200.0 18827.68 -1.16 1500.27 18802.75 -1.42 19073.22
LC 5 5794.88 5.78 5.8 5478.02 0.00 88.7 5477.99 0.00 7200.0 5478.03 0.00 428.7 5477.99 0.00 5477.99
10 8213.51 9.15 11.0 7548.02 0.31 130.0 7525.02 0.00 7200.0 7526.81 0.03 725.4 7525.61 0.01 7524.83
15 8469.11 7.79 27.2 7891.16 0.44 220.5 7863.14 0.08 7200.0 7869.35 0.16 897.9 7858.85 0.03 7856.69
20 10123.49 6.94 59.1 9558.74 0.97 641.7 9506.25 0.42 7200.0 9488.26 0.23 1260.5 9476.98 0.11 9466.51
25 10648.25 5.86 108.4 10227.55 1.68 723.7 10104.99 0.46 7200.0 10088.21 0.29 1525.1 10064.89 0.06 10058.85
30 10394.63 4.84 153.9 10222.89 3.10 619.1 9971.14 0.57 7200.0 9933.83 0.19 1735.6 9901.21 -0.14 9915.06
35 10915.93 0.79 7200.0 10843.66 0.12 1925.4 10817.43 -0.12 10830.43
40 11002.22 0.27 7200.0 10991.20 0.17 1985.1 10951.25 -0.20 10972.65
45 11579.88 0.46 7200.0 11544.00 0.15 2025.8 11510.08 -0.15 11527.10
50 13511.62 0.26 7200.0 13505.18 0.21 2262.0 13454.90 -0.16 13476.69
Avg 9745.82 0.36 7200.0 9726.85 0.17 1477.2 9703.92 -0.07 9710.68
6 LC 50 14336.56 4.96 394.5 14455.92 5.84 2589.5 13694.50 0.26 7200.0 13670.93 0.09 4512.1 13619.39 -0.29 13658.58
100 17630.85 4.50 1823.2 18421.73 9.18 2112.0 16890.54 0.11 7200.0 16805.40 -0.40 6391.5 16736.78 -0.80 16872.38
200 25142.35 4.12 8310.1 27786.21 15.07 828.3 24146.93 0.00 7200.0 24216.63 0.29 7140.9 24063.84 -0.34 24146.93
Avg 19036.59 4.45 3509.3 20221.29 10.95 1843.3 18243.99 0.10 7200.0 18230.99 0.03 6014.8 18140.00 -0.47 18225.96
HC 50 31283.24 2.15 642.5 31412.44 2.57 3081.1 30663.33 0.13 7200.0 30649.56 0.08 4323.8 30581.14 -0.14 30624.92
100 52530.23 1.30 2968.6 53591.81 3.35 2360.4 51865.07 0.02 7200.0 51888.67 0.06 5531.8 51740.12 -0.22 51856.76
200 96076.10 0.82 16435.8 99299.77 4.20 1507.6 95298.01 0.00 7200.0 95576.34 0.29 6927.1 95358.71 0.06 95298.01
Avg 59963.19 1.19 6682.3 61434.67 3.67 2316.4 59275.47 0.03 7200.0 59371.52 0.19 5594.2 59226.66 -0.06 59259.90
Avg Gap 2.84 2.66 0.17 -0.15 -0.38
@@
Results are summarized in Table 1, which reports average values for instances sharing the same values of $n$, $H$, and $C$, on the benchmark instances, small and large (the latter corresponding to the last two blocks). For each method, we report the solution value (Sol), the percentage gap (Gap) relative to the best known solution (BKS, shown in the last column), and the computational time in seconds (T(s)). The gap is computed as $\left(\frac{\mathrm{Sol}}{\mathrm{BKS}}-1\right)\times 100\%$. To the best of our knowledge, the BKS values are compiled from all existing methods for the IRP under an ML-based inventory policy, including the best-performing solutions reported in the DIMACS 2022 Vehicle Routing Implementation Challenge. In contrast to these single-run results, we report the average results of our method over multiple runs, as this provides a more reliable performance estimator given the random nature of the algorithm. Specifically, we present the average solution value (Avg), average gap ($\mathrm{Gap}_{\mathrm{A}}$) to the best known solution (BKS), average CPU time (T(s)), the best solution (Best) among the 10 runs and its corresponding gap to the BKS ($\mathrm{Gap}_{\mathrm{B}}$).
The proposed HGS-IRP algorithm exhibits strong performance in both solution quality and computational efficiency. It consistently produces solutions that are comparable to, or better than, those obtained by existing methods, all within reasonable computational times. Across all 1,040 tested instances, the maximum optimality gap relative to the BKS is only 0.29%, while the average gap is $0.15\%$, both of which outperform all benchmark algorithms. In terms of computational time, HGS-IRP is particularly efficient on large-scale instances, and it also maintains strong solution quality on smaller instances. As the problem size increases, both in number of periods ($H$) and customers ($n$), the advantages of HGS-IRP become more pronounced. On the small 3-day dataset (400 instances), HGS-IRP identifies 37 new best-known solutions, matches the existing BKS in 320 instances, and performs slightly worse in only 43 instances. For the small 6-day dataset, it finds 150 new best solutions and matches the BKS in 83 instances. Most notably, on the large 6-day dataset, HGS-IRP finds new best solutions in 178 out of 240 instances, representing a substantial improvement over the current state-of-the-art in solving large-scale IRP problems.
@@
Experiments on SAHS instances. We further evaluate HGS-IRP on the challenging IRP instances recently introduced by skaalnes2024new. Table 2 presents the average results across instances of varying sizes. skaalnes2024new propose three approaches, and we report the corresponding results: (1) Construction heuristic, (2) Matheuristic, (3) Branch-and-Cut. Only the Branch-and-Cut method provides lower bounds (LB), in contrast to the constructive and matheuristic methods, which are primarily designed to generate feasible solutions and, therefore, only provide upper bounds (UB). Finally, BKS reports the best of the feasible solutions found by the three approaches (the minimum value in UB).
@@
The computational results demonstrate that, out of a total of 270 benchmark instances, the best-known solution is improved in 191 instances and matched in an additional 25 instances. Notably, within the same 2-hour time limit, substantially better solutions are obtained, yielding an average improvement of 18.21% over all instances compared to the BKS previously reported by skaalnes2024new.
Table 2: Comparative analysis on large-scale IRP instances
@@
$n$ UB Gap(%) T(s) UB Gap(%) T(s) LB UB Gap(%) T(s) Avg $\mathrm{Gap}_{\mathrm{A}}$(%) T(s) Best $\mathrm{Gap}_{\mathrm{B}}$(%) BKS
10 7502.74 40.08 0.0 5385.52 0.55 1748.8 5331.14 5356.03 0.00 5422.3 5369.76 0.26 4378.4 5356.81 0.01 5356.03
25 19431.74 40.37 0.0 15072.20 8.88 1800.0 13626.19 13842.88 0.00 7199.1 13886.24 0.31 7078.1 13819.07 -0.17 13842.88
50 42969.72 6.84 0.1 - - 2400.0 28862.31 51389.61 27.78 7193.2 29945.03 -25.54 7390.6 29753.66 -26.02 40218.66
100 86098.23 0.38 0.7 - - 2400.0 52307.74 172930.13 101.61 7193.8 58382.45 -31.94 7215.3 58121.23 -32.24 85775.08
200 154146.38 0.00 6.7 - - 2400.0 76747.91 - - 7217.9 106730.50 -30.76 11139.3 105188.88 -31.76 154146.38